# PINTO: Physics-informed transformer neural operator for learning generalized solutions of partial differential equations for any initial and boundary condition

### v3

@@ Abstract
Applications in physics, engineering, mechanics, and fluid dynamics necessitate solving nonlinear partial differential equations (PDEs) with different initial and boundary conditions. Operator learning, an emerging field, solves these PDEs by employing neural networks to map the infinite-dimensional input and output function spaces. These neural operators are trained using data (observations or simulations) and PDE residuals (physics loss). A key limitation of current neural methods is the need to retrain for new initial/boundary conditions and the substantial simulation data required for training. We introduce a physics-informed transformer neural operator (named PINTO) that generalizes efficiently to new conditions, trained solely with physics loss in a simulation-free setting. Our core innovation is the development of iterative kernel integral operator units that use cross-attention to transform domain points of PDE solutions into initial/boundary condition-aware representation vectors, supporting efficient and generalizable learning. PINTO is utilized to simulate important equations used in fluid mechanics, physics and engineering applications: advection, Burgers, and steady and unsteady Navier-Stokes equations (three flow scenarios). We show that under challenging unseen conditions, the relative errors compared to analytical or numerical (finite difference and volume) solutions are low, merely 20% to 33% of those obtained by other leading physics-informed neural operator methods. Furthermore, PINTO accurately solves advection and Burgers equations at time steps not present in the training points, an ability absent for other neural operators. The code is accessible at https://github.com/quest-lab-iisc/PINTO.
@@ 1 Introduction
PGNN, PINN and PENN learn the map for one instance of the PDE and require re-training for solving with new initial/boundary conditions, much like traditional numerical solvers. On the other hand, neural operators can learn a map between an initial/boundary function and a solution function, thereby achieving generalization to unseen conditions for a particular PDE. However, many of the current leading neural operators such as the Fourier neural operator (FNO) [6, 21, 22, 23, 24, 25, 26, 27, 28, 29], DeepONet [30, 7, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], Graph neural operator [5, 42], Convolutional neural operators [43], Complex neural operator [44], Wavelet neural operator [45, 46, 47], Laplacian neural operator [48, 49], RiemannONets [50], Geometry informed neural operator (GINO) [51], Diffeomorphism Neural Operator [52], Spectral neural operator [53, 54], operator transformer [55, 56, 57], Local neural operator [58], and Peridynamic neural operators [59] are all trained using a vast amount of simulation data. They resemble the PGNN framework; however, due to the fundamental difference in the input-output spaces, neural operators are generalizable.
Recently, physics-informed variants of neural operators (effectively combining the two major categories of neural models to solve PDEs) have been proposed [16, 60, 61]. These methods also need some amount of simulation data (perhaps in a coarse domain) for training. Moreover, as we show in this paper, the current leading architectures that can be trained without any simulation data are not efficient in practice for initial/boundary condition generalization.
@@ 1.1 Related work:
Neural operators to learn maps between infinite-dimensional input and output spaces have gained tremendous popularity in recent years. Concurrent work on branch-trunk network DeepONet architecture, discretization-dependent convolutional neural operators, and discretization-invariant Fourier neural operators has led the way. In the DeepONet architecture, a trunk net encodes the coordinates of the space-time domain of the PDE, and a branch net encodes the discrete input function space [7]. The representations are merged using a Hadamard product. This operator was trained using simulation data. Later work extended the setup to situations where data availability is sparse by introducing PDE-residual loss as a physics-informed regularizer during DeepONet training [30, 71, 36]. Extensions have been reported to multiple input functions [72], efficient numerics [73], and training using the variational form of PDE residuals [35].
@@
Recently, several approaches have been proposed to solve PDEs without using simultation data. For example, a finite element-based physics-informed operator learning framework has been developed with a Galerkin discretized weak formulation and an implicit Euler time integration scheme for temporal discretization as a loss function to incorporate physics into the network [16]. Similarly, the Lippmann-Schwinger operator in Fourier space has been used to construct physical constraints, effectively eliminating Automatic Differentiation [60]. The prior knowledge of the governing equations has been encoded through a stochastic projection-based gradient [77] to solve PDEs for multiple initial and boundary conditions without retraining the model. Still, these approaches come with the cost of approximation and randomness in evaluating gradients.
@@ 1.2 Key contributions
Our main contributions are: (i) development of a new cross-attention mechanism, grounded in neural operator theory, for an efficient neural operator that generalizes the solution of initial boundary value problems for unseen initial/boundary conditions; (ii) training of the neural operator uses only physics loss, without data from numerical simulations; (iii) better simulation than leading physics informed neural operator models for challenging benchmark test cases including the advection equation, Burger’s equation, and Navier Stokes Equations in three different test cases; and (iv) ability to simulate solutions outside the temporal domain used for training. The success of our new neural operator is attributable to the transformation of the PDE solution’s domain query point (spatio-temporal coordinates) into a boundary-aware vector through multiple cross-attention units (iterative kernel integral operator).
@@
$\mathcal{W}_{t}$ Local linear operator
$\mathcal{K}_{t}$ Nonlinear iterative kernel integral operator
$\gamma_{t}$ Bias term
@@
$\mu_{t}$ Hidden representation
$D_{t}$ Domain of hidden representation
@@ 2.2 Cross Attention Neural Operator Theory
where $\mathcal{P}$ is the lifting operator, $\mathcal{Q}$ is the projection operator and the other terms are the $t=1,\cdots,T$ hidden layers that perform iterative kernel integration. Here, $\sigma$ is the pointwise nonlinear activation function, $\mathcal{W}_{t}$ is a local linear operator, $\mathcal{K}_{t}$ is the nonlinear kernel integral operator and $\gamma_{t}$ is the bias function. These layers map each hidden representation to the next, i.e., $\{\mu_{t}\colon D_{t}\to\mathbb{R}^{d_{\mu_{t}}}\}\mapsto\{\mu_{t+1}\colon D_{t+1}\to\mathbb{R}^{d_{\mu_{t+1}}}\}$ using
@@
We propose an encoding layer $\mathcal{P}=MLP_{qpe}(X)$ for the lifting operator and a multi-layer perceptron $\mathcal{Q}=MLP(\mu_{T}(X))$ for the projection operator. For $\mathcal{K}_{t}$, we propose a cross-attention kernel integral operator
$\mathcal{K}_{t}=\int_{\partial\Omega}\frac{\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X),B_{h}MLP_{bpe}\left(X_{b}\right)\rangle}{\sqrt{m}}\right)}{\int_{\partial\Omega}\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X),B_{h}MLP_{bpe}\left(X_{b}\right)\rangle}{\sqrt{m}}\right)\rm{d}X_{b}}R_{h}MLP_{bve}\left(b\right)\,\rm{d}X_{b}\,,$ (6)
@@ 2.3 Practical Implementation
In the second stage, the network obtains a boundary-aware query point encoding vector through multiple passes of our new cross-attention units. In each cross-attention unit, the boundary key and value are shared so that the initial/boundary conditions influence the context-aware hidden representation of query points correctly. The vector representation from the QPE is used as the query of the first cross-attention block, and the output of the previous cross-attention unit is used for all subsequent cross-attention units. We implement a deep multihead cross attention unit to increase the generalization ability and learning capacity of the neural operator model. The discrete multihead cross-attention unit is implemented as follows. First, the attention score $\zeta_{i}$ between the encoded query point at the layer $t$ and every $i$th discrete initial/boundary point and value is calculated as
$\zeta_{i}=\left(\sum_{l=1}^{L}\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X)\;,\;B_{h}MLP_{bpe}(X_{b}^{l})\rangle}{\sqrt{m}}\right)\,\right)^{-1}\text{exp}\left(\frac{\langle A_{h}\mu_{t}(X),\,B_{h}MLP_{bpe}(X_{b}^{i})\rangle}{\sqrt{m}}\right)\,.$ (7)
@@
$\mu_{t+1}(X)=\sigma\left(A_{h}\mu_{t}(X)+\left(\sum_{h=1}^{H}\sum_{i=1}^{L}\zeta_{i}\cdot R_{h}MLP_{bve}(\mathcal{B}(X_{b}^{i}))\right)\right)\,.$ (8)
@@ 3 Applications of PINTO
The performance of the PINTO model for all numerical examples is compared to the physics-informed DeepONet model (PI-DeepONet) [7, 36]. Although PI-DeepONet has not been employed to learn generalized PDE solutions for unseen initial/boundary conditions akin to our problem, we adapted it to obtain PDE solutions under various initial/boundary conditions by treating the input functions for the branch network as initial and boundary conditions. PI-DeepONet is chosen as the benchmark for comparison as it can be trained using only physics-loss unlike other neural operators that need some amount of simulation data. Table 2 summarizes the relative error for all test cases. The definition of relative error and the results of other metrics to quantify the error are provided in  A. In the upcoming sections, we take a deep dive into each test case. The code and data for all test cases are available on Zenodo [78] and GitHub [^1] https://github.com/quest-lab-iisc/PINTO.
@@
$1$D Burgers equation $4.81\%\,\left(4.43\%\right)$ $5.24\%\,\left(4.51\%\right)$ $12.81\%\,\left(11.85\%\right)$ $15.03\%\,\left(10.78\%\right)$
@@ 4 Discussion
Limitations: Since the training of the PINTO model uses collocation points and physics-loss, it may suffer spectral bias, a common phenomenon while training neural networks, and has also been reported in other PINN models. There might also be training discrepancies due to imbalances in different terms in the loss functions. Adaptive training, adaptive sampling, and dynamic weighting strategies [88, 89, 90, 91, 92, 93, 94] must be adopted to address these limitations.
@@ Appendix A Performance metrics
Kovasznay $0.0007$ $0.0006$ $0.0114$ $0.0035$ $0.0017$ $0.00151$ $0.0573$ $0.0398$
Beltrami $0.0295$ $0.0171$ $0.0336$ $0.0227$ $0.1276$ $0.0848$ $0.2709$ $0.1529$
Lid Driven $0.0527$ $0.0494$ $0.077$ $0.073$ $0.0905$ $0.0739$ $0.2189$ $0.1708$

### v4

@@ Abstract
Applications in physics, engineering, mechanics, and fluid dynamics necessitate solving nonlinear partial differential equations (PDEs) with different initial and boundary conditions. Operator learning, an emerging field, solves these PDEs by employing neural networks to map the infinite-dimensional input and output function spaces. These neural operators are trained using data (observations or simulations) and PDE residuals (physics loss). A key limitation of current neural methods is the need to retrain for new initial/boundary conditions and the substantial simulation data required for training. We introduce a physics-informed transformer neural operator (named PINTO) that generalizes efficiently to new conditions, trained solely with physics loss in a simulation-free setting. Our core innovation is the development of iterative kernel integral operator units that use cross-attention to transform domain points of PDE solutions into initial/boundary condition-aware representation vectors, supporting efficient and generalizable learning. The working of PINTO is demonstrated by simulating important 1D and 2D equations used in fluid mechanics, physics and engineering applications: advection, Burgers, and steady and unsteady Navier-Stokes equations (three flow scenarios). We show that under challenging unseen conditions, the relative errors compared to analytical or numerical (finite difference and volume) solutions are low, merely 20% to 33% of those obtained by other leading physics-informed neural operator methods. Furthermore, PINTO accurately solves advection and Burgers equations at time steps not present in the training points, an ability absent for other neural operators. The code is accessible at https://github.com/quest-lab-iisc/PINTO.
@@ 1 Introduction
PGNN, PINN and PENN learn the map for one instance of the PDE and require re-training for solving with new initial/boundary conditions, much like traditional numerical solvers. On the other hand, neural operators can learn a map between an initial/boundary function and a solution function, thereby achieving generalization to unseen conditions for a particular PDE. However, many of the current leading neural operators such as the Fourier neural operator (FNO) [6, 21, 22, 23, 24, 25, 26, 27, 28, 29], DeepONet [30, 7, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], Graph neural operator [5, 43], Convolutional neural operators [44], Complex neural operator [45], Wavelet neural operator [46, 47, 48], Laplacian neural operator [49, 50], RiemannONets [51], Geometry informed neural operator (GINO) [52], Diffeomorphism Neural Operator [53], Spectral neural operator [54, 55], operator transformer [56, 57, 58], Local neural operator [59], and Peridynamic neural operators [60] are all trained using a small to large amount of simulation data. They resemble the PGNN framework; however, due to the fundamental difference in the input-output spaces, neural operators are generalizable. Some of the above neural operators have a physics-informed variant [61, 30]. However, these also require some amount of simulation data (perhaps in a coarse domain) for training. Moreover, as we show in this paper, the current leading architectures that can be trained without any simulation data are not efficient in practice for initial/boundary condition generalization.
@@ 1.1 Related work:
Neural operators to learn maps between infinite-dimensional input and output spaces have gained tremendous popularity in recent years. Concurrent work on branch-trunk network DeepONet architecture, discretization-dependent convolutional neural operators, and discretization-invariant Fourier neural operators has led the way. In the DeepONet architecture, a trunk net encodes the coordinates of the space-time domain of the PDE, and a branch net encodes the discrete input function space [7]. The representations are merged using a dot product. This operator was trained using simulation data. Later work extended the setup to situations where data availability is sparse by introducing PDE-residual loss as a physics-informed regularizer during DeepONet training [30, 36]. Extensions have been reported to multiple input functions [71], efficient numerics [72], and training using the variational form of PDE residuals [35].
@@
Recently, several approaches have been proposed to solve PDEs without using simulation data. For example, a finite element-based physics-informed operator learning framework has been developed with a Galerkin discretized weak formulation and an implicit Euler time integration scheme for temporal discretization as a loss function to incorporate physics into the network [16]. Similarly, the Lippmann-Schwinger operator in Fourier space has been used to construct physical constraints, effectively eliminating Automatic Differentiation [75]. The prior knowledge of the governing equations has been encoded through a stochastic projection-based gradient [76] to solve PDEs for multiple initial and boundary conditions without retraining the model. Still, these approaches come with the cost of approximation and randomness in evaluating gradients.
@@ 1.2 Key contributions
Our main contributions are: (i) development of a new cross-attention mechanism, grounded in neural operator theory, for an efficient neural operator that generalizes the solution of initial boundary value problems for unseen initial/boundary conditions; (ii) training of the neural operator uses only physics loss, without data from numerical simulations; (iii) better simulation than leading physics informed neural operator models for challenging benchmark 1D and 2D test cases including the advection equation, Burger’s equation, and Navier Stokes Equations in three different test cases; and (iv) ability to simulate solutions outside the temporal domain used for training. The success of our new neural operator is attributable to the transformation of the PDE solution’s domain query point (spatio-temporal coordinates) into a boundary-aware vector through multiple cross-attention units (iterative kernel integral operator).
@@
$\mathcal{W}_{j}$ Local linear operator
$\mathcal{K}_{j}$ Nonlinear iterative kernel integral operator
$\gamma_{j}$ Bias term
@@
$\mu_{j}$ Hidden representation
$D_{j}$ Domain of hidden representation
@@ 2.2 Cross Attention Neural Operator Theory
where $\mathcal{P}$ is the lifting operator, $\mathcal{Q}$ is the projection operator and the other terms are the $j=1,\cdots,T$ hidden layers that perform iterative kernel integration. Here, $\sigma$ is the pointwise nonlinear activation function, $\mathcal{W}_{j}$ is a local linear operator, $\mathcal{K}_{j}$ is the nonlinear kernel integral operator and $\gamma_{j}$ is the bias function. These layers map each hidden representation to the next, i.e., $\{\mu_{j}\colon D_{j}\to\mathbb{R}^{d_{\mu_{j}}}\}\mapsto\{\mu_{j+1}\colon D_{j+1}\to\mathbb{R}^{d_{\mu_{j+1}}}\}$ using
@@
We propose an encoding layer $\mathcal{P}=MLP_{qpe}(X)$ for the lifting operator and a multi-layer perceptron $\mathcal{Q}=MLP(\mu_{T}(X))$ for the projection operator. For $\mathcal{K}_{j}$, we propose a cross-attention kernel integral operator
$\mathcal{K}_{j}=\int_{\partial\Omega}\frac{\text{exp}\left(\frac{\langle A_{h}\mu_{j}(X),B_{h}MLP_{bpe}\left(X_{b}\right)\rangle}{\sqrt{m}}\right)}{\int_{\partial\Omega}\text{exp}\left(\frac{\langle A_{h}\mu_{j}(X),B_{h}MLP_{bpe}\left(X_{b}\right)\rangle}{\sqrt{m}}\right)\rm{d}X_{b}}R_{h}MLP_{bve}\left(b\right)\,\rm{d}X_{b}\,,$ (6)
@@ 2.3 Practical Implementation
In the second stage, the network obtains a boundary-aware query point encoding vector through multiple passes of our new cross-attention units. In each cross-attention unit, the boundary key and value are shared so that the initial/boundary conditions influence the context-aware hidden representation of query points correctly. The vector representation from the QPE is used as the query of the first cross-attention block, and the output of the previous cross-attention unit is used for all subsequent cross-attention units. We implement a deep multihead cross attention unit to increase the generalization ability and learning capacity of the neural operator model. The discrete multihead cross-attention unit is implemented as follows. First, the attention score $\zeta_{i}$ between the encoded query point at the layer $j$ and every $i$th discrete initial/boundary point and value is calculated as
$\zeta_{i}=\left(\sum_{l=1}^{L}\text{exp}\left(\frac{\langle A_{h}\mu_{j}(X)\;,\;B_{h}MLP_{bpe}(X_{b}^{l})\rangle}{\sqrt{m}}\right)\,\right)^{-1}\text{exp}\left(\frac{\langle A_{h}\mu_{j}(X),\,B_{h}MLP_{bpe}(X_{b}^{i})\rangle}{\sqrt{m}}\right)\,.$ (7)
@@
$\mu_{j+1}(X)=\sigma\left(A_{h}\mu_{j}(X)+\left(\sum_{h=1}^{H}\sum_{i=1}^{L}\zeta_{i}\cdot R_{h}MLP_{bve}(\mathcal{B}(X_{b}^{i}))\right)\right)\,.$ (8)
@@ 3 Applications of PINTO
The performance of the PINTO model for all numerical examples is compared to the physics-informed DeepONet model (PI-DeepONet) [7, 36]. PI-DeepONet is chosen as the benchmark for comparison as it can be trained using only physics-loss unlike other neural operators that need some amount of simulation data. Table 2 summarizes the relative error for all test cases. The definition of relative error and the results of other metrics to quantify the error are provided in  A. In the upcoming sections, we take a deep dive into each test case. The code and data for all test cases are available on Zenodo [77] and GitHub [^1] https://github.com/quest-lab-iisc/PINTO.
@@
$1$D Burgers equation $4.81\%\,\left(4.43\%\right)$ $5.24\%\,\left(4.51\%\right)$ $12.81\%\,\left(15.03\%\right)$ $11.85\%\,\left(10.78\%\right)$
@@ 4 Discussion
Limitations: Since the training of the PINTO model uses collocation points and physics-loss, it may suffer spectral bias, a common phenomenon while training neural networks, and has also been reported in other PINN models. There might also be training discrepancies due to imbalances in different terms in the loss functions. Adaptive training, adaptive sampling, and dynamic weighting strategies [87, 88, 89, 90, 91, 92, 93, 94] must be adopted to address these limitations.
@@ Appendix A Performance metrics
Kovasznay $0.000739$ $0.000662$ $0.000715$ $0.000655$ $0.00179$ $0.00151$ $0.001758$ $0.00149$
Beltrami $0.0295$ $0.0171$ $0.0249$ $0.0130$ $0.1276$ $0.0848$ $0.1046$ $0.065$
Lid Driven $0.0522$ $0.0494$ $0.0817$ $0.073$ $0.0874$ $0.0739$ $0.2523$ $0.1708$